OpenStudy (anonymous):

what is the indefinite integral of tan(x)^4dx I believe I have the answer but I'm not sure

OpenStudy (lgbasallote):

\(\int (\tan x)^4 dx\) right? not \(\int \tan (x^4)dx\)

OpenStudy (anonymous):

yep tan(x) to the 4th

OpenStudy (anonymous):

Give me a second to type it up.

OpenStudy (anonymous):

i just want to know really if the answer is what I got ((tan(x)^(3)/3)-(tan(x))+(x)+C)

OpenStudy (lgbasallote):

we can use trigonometric substitution \(\int \tan^2 x (\tan^2 x)dx\) then let tan^2 x = sec^2 x - 1 \(\int \tan^2x (\sec^2x - 1) dx\) now distribute... \(\int (\tan^2 x \sec^2 x)dx - \int \tan^2 xdx\) \(\int (\tan^2 x \sec^2 x)dx - \int (\sec^2 x - 1)dx\) \(\int (\tan^2 x \sec^2 x)dx - \int \sec^2 xdx - \int dx\) you an do it from tehre :p

OpenStudy (anonymous):

I got that ^^^^

OpenStudy (anonymous):

(1/3)tan^3(x)+x-tan(x)+C

OpenStudy (lgbasallote):

oh oh yeah should be + int dx

OpenStudy (anonymous):

Ok thanks my professor doesn't give the answers to exams if you get them wrong so I re did it and I guess I did it right so thanks guys